# Scaling Environments for Organoid Intelligence with LLM-Automated Design and Plasticity-Based Evaluation

### v1

@@ A.2.1 Scaling the conditional avoidance environment
# Shift the boundary occasionally, making the task harder
if random.choice(range(0, 1 )) == 1:
@@ A.2.2 Scaling the predator-prey environment
# Encode prey position using frequency on group C
@@
# Encode predator position on group D
@@
# Use different frequency/amplitude to encode adversary
@@ A.2.3 Scaling the dynamic interception (Pong) environment
# 3. Action decoding (paddle movement)
@@
# Each organoid receives the same information
@@
# Optional: deliver small reward to hitting player
@@ A.3.1 Example 1: JSON-based parameter optimization
This approach is ideal for systematic exploration of a known parameter space. By constraining the LLM to output a JSON object, we ensure the protocol is always syntactically correct and can be easily parsed and validated before execution. This method is particularly useful for optimizing feedback mechanisms, such as reward and punishment, by exploring different stimulation modalities and their specific parameters.
The following prompt (Listing 13) asks the LLM to propose a new set of parameters for the predator-prey task, taking into account historical data and a new hypothesis. It must decide between an electrical reward and a chemical (dopamine uncaging) reward, and then specify the nested parameters for its choice.
@@
Hypothesis: A stronger, more salient electrical reward might be more effective than the previous attempts. Lets try a tri-phasic pulse to ensure charge balance and a higher amplitude.
@@
An example of an LLM response from our trials is shown in Listing 14. The model correctly interprets the hypothesis, selects the electrical modality, and populates the nested parameters accordingly, including the requested tri-phasic shape and a significantly higher amplitude_uA. This structured output can be directly loaded by the experimental control software to configure the next run.
@@ A.3.2 Example 2: Full script generation for curriculum design
This approach grants the LLM maximum flexibility to design novel experimental logic, making it suitable for tasks like curriculum design, where the protocol itself must change based on agent performance. By providing the LLM with a detailed API specification, it can compose complex sequences of operations that would be impossible to specify in a simple key-value format.
The prompt in Listing 15 outlines a more comprehensive API and asks the LLM to write a complete Python script for a block of trials. This includes defining the logic for state encoding, action decoding, feedback delivery, and, crucially, a terminal condition with a suggestion for the next curriculum step. This example uses FinalSpark’s Neuroplatform [Jordan et al., 2024]
@@
The resulting script (Listing 16) is an example from our trials. It demonstrates the LLM’s ability to correctly use the specified API. It correctly structures the trial loop, implements the state encoding and action decoding logic, and calls the complex fire_pulse_train function with the correct parameters for both reward and punishment. Finally, it calculates the performance metric and prints a context-aware suggestion, effectively automating a step in curriculum design.
@@
if predator_pos != prey_pos: # Punishment if not captured
@@
electrodes=SENSORY_ELECTRODES, shape=’bi-phasic’,
@@
print("Suggestionfornextblock:Lowperformance.Repeatblock.")

### v2

@@ A.2.1 Scaling the conditional avoidance environment
# Shift the boundary occasionally, forcing the agent’s world model to adapt
if random.choice([True, False]):
@@ A.2.2 Scaling the predator-prey environment
# Encode prey position using spatial location on group C and frequency for y-axis
@@
# Encode predator position similarly on group D
@@
# Use different frequency/amplitude to encode adversary on a sensory channel
@@ A.2.3 Scaling the dynamic interception (Pong) environment
# 3. Action decoding (paddle movement, now horizontal)
@@
# Each organoid receives the same world state information
@@
# Optional: deliver small reward to hitting player to encourage long rallies
@@ A.3.1 Example 1: JSON-based parameter optimization
This approach is ideal for systematic exploration of a known parameter space. Constraining the LLM to output a JSON object ensures the protocol is syntactically correct and can be easily parsed and validated. This method is useful for optimizing feedback mechanisms by exploring different stimulation modalities and their parameters. The prompt (Listing 13) asks the LLM to propose new parameters for the predator-prey task, taking into account historical data and a new hypothesis.
@@
Hypothesis: A stronger, more salient electrical reward might be more effective. Lets try a tri-phasic pulse to ensure charge balance and a higher amplitude.
@@
An example of an LLM response is shown in Listing 14. The model correctly interprets the hypothesis, selects the electrical modality, and populates the nested parameters accordingly, including the requested tri-phasic shape and a significantly higher amplitude. This structured output can be directly loaded by the experimental control software.
@@ A.3.2 Example 2: Full script generation for curriculum design
This approach grants the LLM maximum flexibility to design novel experimental logic, suitable for tasks like curriculum design where the protocol itself must change based on agent performance. By providing the LLM with a detailed API specification, it can compose complex sequences of operations. The prompt in Listing 15 asks the LLM to write a complete Python script for a block of trials for a hypothetical platform. This includes defining the logic for state encoding, action decoding, feedback delivery, and a terminal condition with a suggestion for the next curriculum step. This example uses FinalSpark’s Neuroplatform [Jordan et al., 2024].
@@
The resulting script (Listing 16) demonstrates the LLM’s ability to correctly use the specified API. It structures the trial loop, implements the state encoding and action decoding logic, calls the ‘fire_pulse_train‘ function with correct parameters for reward and punishment, and prints a context-aware suggestion, effectively automating a step in curriculum design.
@@
if predator_pos != prey_pos: # Punishment if not captured in time
@@
electrodes=random.sample(SENSORY_ELECTRODES, 4), shape=’bi-phasic’,
@@
print("Suggestionfornextblock:Lowperformance.Repeatblockwithincreasedrewardsaliency.")